For a solution of pH 5.0 at 25°C, calculate [H+] and [OH-].


1
Expert's answer
2022-02-23T23:14:02-0500

We know that pH+pOH=14
So, pOH=14−pH=14−5=9
Also, pOH=−log [OH−]=9
Hence, [OH−]=1×10−9 M
Therefore, the hydroxide ion concentration of the solution is 10−9 M.